# Revising Multimodal VAEs with Diffusion Decoders

### v1

@@ Abstract
Multimodal VAEs often struggle with generating high-quality outputs, a challenge that extends beyond the inherent limitations of the VAE framework. The core issue lies in the restricted joint representation of the latent space, particularly when complex modalities like images are involved. Feedforward decoders, commonly used for these intricate modalities, inadvertently constrain the joint latent space, leading to a degradation in the quality of the other modalities as well. Although recent studies have shown improvement by introducing modality-specific representations, the issue remains significant. In this work, we demonstrate that incorporating a flexible diffusion decoder specifically for the image modality not only enhances the generation quality of the images but also positively impacts the performance of the other modalities that rely on feedforward decoders. This approach addresses the limitations imposed by conventional joint representations and opens up new possibilities for improving multimodal generation tasks using the multimodal VAE framework. Our model provides state-of-the-art results compared to other multimodal VAEs in different datasets with higher coherence and superior quality in the generated modalities
@@ 1 Introduction
Deep learning has revolutionized many fields through its ability to extract and learn patterns that are otherwise difficult through other means (LeCun et al., 2015; Schmidhuber, 2015). One area that has seen immense progress is representation learning, where deep neural networks learn rich representations that disentangle the underlying explanatory factors of data by projecting the data into a latent space (Bengio et al., 2013). Variational autoencoders (VAEs) have emerged as a powerful framework for learning interpretable representations in an unsupervised manner (Kingma and Welling, 2014; Rezende et al., 2014). VAEs were initially built to process a single modality, such as an image. However, in the real world, data often comes from multiple modalities that are interconnected and expressed in different formats (Baltrusaitis et al., 2019). Multimodal learning aims to build models that can process and relate information from various modalities, such as vision, language, and other modalities together (Ngiam et al., 2011). Incorporating this multimodal information can have a huge impact on the performance of the model on tasks that require understanding from multiple perspectives (Srivastava and Salakhutdinov, 2014).
@@
Multimodal variational autoencoders extend the VAE framework to learn joint representations across modalities (Ngiam et al., 2011; Suzuki et al., 2017; Wu and Goodman, 2018). By capturing the correlations between modalities, these models can learn robust and informative representations. These learned multimodal representations can be used in different ways, including transfer learning in downstream tasks (Hwang et al., 2021; Jing et al., 2024), and they also provide a framework to perform unconditional and conditional generation across the different modalities (Shi et al., 2019; Wesego and Rooshenas, 2023).
@@
Figure 1: Architecture of our model where modality-specific encoders map the input data from different modalities into latent space which are then fused using PoE and passed to the respective decoders.
@@
Figure 2: Unconditional sampling technique that leverages a score-based model to transform a Gaussian noise distribution into the PoE posterior distribution.
@@
Despite their promising potential, multimodal VAEs continue to face significant challenges in effectively integrating diverse modalities, generating high-quality samples, and scaling to high-dimensional data (Daunhawer et al., 2022; Wesego and Rooshenas, 2023). Various prior works have attempted to address these issues by proposing alternative ELBO objectives, yet the core problems still persist in multimodal VAEs (Sutter et al., 2021, 2020). Due to these ongoing limitations, most multimodal studies have been constrained to low-dimensional datasets such as MNIST or PolyMNIST (LeCun and Cortes, 2010; Sutter et al., 2020). Additionally, Daunhawer et al. (2022) highlights the generative discrepancies in mixture-based models and raises concerns about the practicality of these models in real-world applications.
@@
On the other hand, diffusion models have emerged as a powerful class of generative models, gaining substantial attention in recent years for their ability to produce high-quality samples Ho et al. (2020); Dhariwal and Nichol (2021). These models operate by gradually transforming a data distribution into a noise distribution, with the goal of learning to reverse this process. The forward process is a Markovian sequence that incrementally adds noise to the data distribution over several timesteps, while the reverse process, typically modeled by a neural network, learns to remove the noise and reconstruct the data (Ho et al., 2020). A notable advantage of diffusion models is their capacity to generate exceptionally high-quality samples, achieving state-of-the-art performance in various generative tasks, including image synthesis and inpainting (Saharia et al., 2022; Lugmayr et al., 2022). Preechakul et al. (2022) introduced diffusion autoencoder, where the encoder resembles that of a VAE, but the decoder is modeled using a diffusion process. The key difference between a diffusion decoder and a traditional feedforward decoder used in VAEs is that the diffusion decoder conditions the generation process on the latent representation rather than directly feeding it as input. This approach enables the diffusion autoencoder to learn meaningful representations without being constrained by the output size.
In this study, we propose a novel multimodal generative model that leverages the strengths of both diffusion autoencoders and multimodal VAEs by learning a unified joint representation across multiple modalities. In traditional multimodal VAEs, joint representation often becomes restrictive, especially when rich representations are required for complex modalities like images, leading to the need for modality-specific representations Palumbo et al. (2023); Hwang et al. (2021). However, we demonstrate that by employing a flexible diffusion decoder, the joint representation alone suffices to generate high-quality outputs not only for complex modalities but also for other modalities using feedforward decoders.
@@
To achieve this, we use diffusion decoders for modalities where feedforward decoders have shown limitations, such as images, ensuring high-quality sample generation and improved representation capacity. Conversely, for modalities where feedforward or transformer-based decoders have proven effective, like text or sparse outputs, we retain standard VAE decoders. This hybrid approach allows our model to harness the strengths of each generative model class while effectively addressing their respective limitations.
@@
Our model achieves better coherence in conditional generation among the generated modalities while generating high-quality samples. We use the product of experts (PoE) to fuse the representations from different modalities (Wu and Goodman, 2018). This approach enables our model to capture the correlations and dependencies between modalities. For unconditional generation tasks, where the goal is to generate samples from all modalities simultaneously without conditioning on any specific modality, we propose the use of an auxiliary model. This auxiliary model transforms a Gaussian noise distribution into the approximate posterior product of experts distribution. By leveraging this technique, our model gains increased flexibility and can perform well in both conditional and unconditional generation settings.
@@
Figure 2 illustrates the general design and architecture of our proposed model. The modality-specific encoders map the input data from different modalities into latent representations, which are then fused using the PoE mechanism. The fused representation is then passed to the respective decoders, which can be either diffusion autoencoders or standard VAEs, depending on the modality. Figure 2 shows how to perform unconditional sampling where the score-based model is employed to provide the necessary input distributions to the decoders.
@@
  • We introduce a novel generative multimodal autoencoder architecture that seamlessly integrates diffusion autoencoders with standard autoencoders. This hybrid design allows our model to generate high-quality samples across modalities without compromising the coherence and consistency among the generated modalities. By doing so, our architecture strikes an optimal balance between sample quality and cross-modal coherence.
  • To enhance the model’s performance in unconditional generation settings, we propose a novel sampling technique that employs an auxiliary score-based model that transforms a Gaussian noise distribution into the approximate distribution of the PoE posterior to improve the model’s ability to generate coherent and diverse samples across all modalities in an unconditional manner.
  • Our approach can be scaled to high-dimensional data modalities with good generative quality compared to previous approaches
Through these contributions, our model addresses the limitations of previous multimodal VAE approaches, while retaining their desirable properties and extending their capabilities to handle high-dimensional data modalities effectively. Our approach paves the way for more powerful and flexible multimodal generative models capable of generating high-quality samples with enhanced coherence.
@@ 2 Related Works
As the path to increase the performance and understanding of neural networks continues, multimodal learning has emerged as a promising research direction in the deep learning field that has started from early works such as Ngiam et al. (2011); Tsai et al. (2019). Ngiam et al. (2011) demonstrated the potential of multimodal deep learning models in leveraging information from multiple modalities to improve performance on multiple tasks by learning combined features. As variational autoencoders (VAEs) gained traction for simultaneously learning representations and generation models within individual modalities, researchers soon explored extending the VAE framework to handle multiple modalities simultaneously, giving rise to multiple variants of multimodal VAEs.
@@
Alternatively, Wesego and Rooshenas (2023); Bounoua et al. (2024) proposed to use unimodal VAEs to represent each modality and use an additional neural network to fuse the modalities instead of a product/mixture of experts using a two-step training process. They both select a score-based network to join the modalities in the latent space and sample from the joint distribution. Rombach et al. (2021) proposed text-to-image generation that employs a high-quality autoencoder to compress images into latent space, allowing for diffusion processes to occur in this reduced dimensionality. They integrated CLIP models using cross-attention to effectively represent and condition text inputs. Building upon this,  Xu et al. (2023); Tang et al. (2023) proposed multi-modal latent diffusion models that can generate any modality from any given condition. They expand Rombach et al. (2021) work to use multiple diffusion flows to achieve this. It’s worth noting that these advanced models require training on extremely large-scale datasets which requires extensive computational resources for the training processes.
@@ 3.1 Diffusion Autoencoder
A diffusion model can be formulated as a latent variable model where the marginal distribution is expressed as $p_{\theta}(\mathbf{x}_{0})=\int p_{\theta}(\mathbf{x}_{0:T})\mathrm{d}\mathbf{x}_{1:T}$ (Ho et al., 2020). The combined probability distribution $p_{\theta}(\mathbf{x}_{0:T})=p(\mathbf{x}_{T})\prod_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ is the reverse joint process where $p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{0},\mathbf{I})$. One can sample any intermediate time step $t$ in the forward process starting from the data point $\mathbf{x}_{0}$ using the Gaussian distribution $q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$ to add noise to the data. A neural network is trained to reverse this process where $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t))$. Ho et al. (2020) observed that using a simplified objective and predicting the noise is better and easier to implement. The objective is shown as follows in equation  1:
@@
A diffusion autoencoder combines the architectural setup of an autoencoder with a diffusion decoder (Preechakul et al., 2022). An encoder $q(\mathbf{z|x})$ projects the data into a latent representation $\mathbf{z}$. This latent representation is fed to the diffusion decoder which generates samples conditioned on $\mathbf{z}$. By training the model this way, the diffusion model can be used as a powerful representation where the encoder learns a meaningful latent representation of the data similar to an autoencoder (Hudson et al., 2023). The objective for a diffusion autoencoder is very similar to the diffusion objective, except we have additional conditioning coming from the encoder. It is trained by minimizing $\mathbb{E}_{\mathbf{x_{0},\epsilon_{t}}}[\left\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},\mathbf{z},t)\right\|^{2}]$ where $\mathbf{z}$ comes from the encoder network $q(\mathbf{z|x})$.
@@ 3.2 Multimodal VAE with Diffusion Decoder
Assuming that we have M modalities, the observed data is represented as $\mathbf{x}={\mathbf{x}^{(1)},\mathbf{x}^{(2)},\ldots,\mathbf{x}^{(M)}}$. A multimodal ELBO objective optimizes the lower bound of the log-likelihood $\log p_{\theta}(\mathbf{x})$. In the multimodal setting, we can write the ELBO as shown in equation  2 where $\mathbf{x}$ represents the multimodal data.
$\begin{split}\log p(\mathbf{x})\geq\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}\left[\log p_{\theta}(\mathbf{x}|\mathbf{z})\right]-D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x})\|p(\mathbf{z}))\end{split}$ (2)
Daunhawer et al. (2022) showed that for some subset $\mathcal{S}$ from the combination of the modalities, a lower bound can be derived, shown in equation 3, where different subsets can give rise to different multimodal VAE objectives of previous works of MoPoE, MMVAE, and MVAE Sutter et al. (2021); Shi et al. (2019); Wu and Goodman (2018).
$\begin{split}\log p(\mathbf{x})\geq&\sum_{A\in\mathcal{S}}\omega_{A}\Big{\{}\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x}_{A})}\left[\log p_{\theta}(\mathbf{x}|\mathbf{z})\right]-D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x}_{A})\|p(\mathbf{z}))\Big{\}}.\end{split}$ (3)
If we assume the modalities are conditionally independent given the latent representation $\mathbf{z}$, we can express it as $p_{\theta}(\mathbf{x_{1}},\mathbf{x_{2}},...,\mathbf{x_{M}})=p({\mathbf{z}})p_{\theta}(\mathbf{x}_{1}|\mathbf{z})p_{\theta}(\mathbf{x}_{2}|\mathbf{z})...p_{\theta}(\mathbf{x}_{M}|\mathbf{z})$. Then, the ELBO can be written as shown in equation 4:
@@
$\begin{split}\log p(\mathbf{x})\geq&\sum_{A\in\mathcal{S}}\omega_{A}\Bigg{\{}\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x}_{A})}\left[\sum_{i}^{M}\log p_{\theta}(\mathbf{x}_{i}|\mathbf{z})\right]-D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x}_{A})\|p(\mathbf{z}))\Bigg{\}}.\end{split}$ (4)
In our case, we want to use diffusion decoders for image modalities as the generative model while we still want to maintain support for standard VAEs for some modalities. In order to achieve this, we can introduce two indicator functions, one for using a diffusion autoencoder and another for using a standard VAE. Let’s define $\mathbb{I}_{\text{diff}^{(m)}}$ as the indicator function for using a diffusion autoencoder for modality $m$, and $\mathbb{I}_{\text{vae}^{(m)}}$ as the indicator function for using a standard VAE for modality $m$. Then, we can express the objective as shown in equation 5.
@@
$\begin{split}\sum_{A\in\mathcal{S}}\omega_{A}\Bigg{\{}&\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x}_{A})}\Bigg{[}\sum_{i}^{M}\mathbb{I}_{\text{vae}^{(i)}}\log p_{\theta}(\mathbf{x}_{i}|\mathbf{z})+\mathbb{I}_{\text{diff}^{(i)}}\mathbb{E}_{\mathbf{x_{i},\epsilon_{t}}}\Big{[}\left\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{i_{t}},\mathbf{z},t)\right\|^{2}\Big{]}\Bigg{]}\\ &-D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x}_{A})\|p(\mathbf{z}))\Bigg{\}}\end{split}$ (5)
The above formulation may not still be a valid lower bound on the log-likelihood, but it provides a seamless way to integrate the diffusion decoder into the multimodal VAE setup. We choose the product of experts as the posterior where one subset is sampled from the powerset of modalities. In addition to this, to increase the alignment between representations of different modalities, we add a KL-regularization term between the posterior and the individual encoder outputs. The final objective is shown in equation 6 where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters set to $1e^{-5}$ because higher values caused instability during training. We call our proposed model, which combines diffusion autoencoders and variational autoencoders for multimodal generative modeling, Multimodal Diffusion Decoder Variational Autoencoder (MDDVAE).
@@
$\begin{split}\sum_{A\in\mathcal{S}}\omega_{A}\Bigg{\{}&\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x}_{A})}\Bigg{[}\sum_{i}^{M}\mathbb{I}_{\text{vae}^{(i)}}\log p_{\theta}(\mathbf{x}_{i}|\mathbf{z})+\mathbb{I}_{\text{diff}^{(i)}}\mathbb{E}_{\mathbf{x_{i},\epsilon_{t}}}\Big{[}\left\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{i_{t}},\mathbf{z},t)\right\|^{2}\Big{]}\Bigg{]}\\ &-\lambda_{1}D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x}_{A})\|p(\mathbf{z}))-\lambda_{2}\sum_{i}^{M}\Big{[}D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x}_{A})\|q(\mathbf{z}|\mathbf{x}_{i}))\Big{]}\Bigg{\}}\end{split}$ (6)

### v2

# Multimodal ELBO with Diffusion Decoders

@@ Abstract
Multimodal variational autoencoders have demonstrated their ability to learn the relationships between different modalities by mapping them into a latent representation. Their design and capacity to perform any-to-any conditional and unconditional generation make them appealing. However, different variants of multimodal VAEs often suffer from generating low-quality output, particularly when complex modalities such as images are involved. In addition to that, they frequently exhibit low coherence among the generated modalities when sampling from the joint distribution. To address these limitations, we propose a new variant of the multimodal VAE ELBO that incorporates a better decoder using a diffusion generative model. The diffusion decoder enables the model to learn complex modalities and generate high-quality outputs. The multimodal model can also seamlessly integrate with a standard feed-forward decoder for different types of modality, facilitating end-to-end training and inference. Furthermore, we introduce an auxiliary score-based model to enhance the unconditional generation capabilities of our proposed approach. This approach addresses the limitations imposed by conventional multimodal VAEs and opens up new possibilities to improve multimodal generation tasks. Our model provides state-of-the-art results compared to other multimodal VAEs in different datasets with higher coherence and superior quality in the generated modalities.
@@ 1 Introduction
Deep learning has revolutionized many fields through its ability to extract and learn patterns that are otherwise difficult through other means (LeCun et al., 2015; Schmidhuber, 2015). One area that has seen immense progress is generative models, where neural networks learn to generate samples that are similar to the training distribution. Variational autoencoders (VAEs) have emerged as one powerful framework for this task (Kingma and Welling, 2014; Rezende et al., 2014). VAEs were initially built to process a single modality, such as an image. However, in the real world, data often comes from multiple modalities that are interconnected and expressed in different formats (Baltrusaitis et al., 2019). Multimodal learning aims to build models that can process and relate information from various modalities, such as vision, language, and other modalities together (Ngiam et al., 2011). Incorporating this multimodal information can have a huge impact on the performance of the model in tasks that require understanding from multiple perspectives (Srivastava and Salakhutdinov, 2014).
@@
Multimodal variational autoencoders extend the VAE framework to learn joint representations across modalities (Ngiam et al., 2011; Suzuki et al., 2017; Wu and Goodman, 2018). By capturing the correlations between modalities, these models can learn robust and informative representations. These learned multimodal representations can be used in different ways, including transfer learning in downstream tasks (Hwang et al., 2021; Jing et al., 2024), and they also provide a framework to perform unconditional and conditional generation across the different modalities (Shi et al., 2019; Wesego and Rooshenas, 2024).
@@
Figure 1: Architecture of our model where modality-specific encoders map the input data from different modalities into latent space, which are then fused using a mixture of PoE and passed to the respective decoders.
@@
Figure 2: Unconditional sampling technique that leverages a score-based model to transform a Gaussian noise distribution into the PoE posterior distribution of all modalities.
@@
Despite their promising potential, multimodal VAEs continue to face significant challenges in effectively integrating diverse modalities, generating high-quality samples, and scaling to high-dimensional data (Daunhawer et al., 2022; Wesego and Rooshenas, 2024). Various prior works have attempted to address these issues by proposing alternative ELBO objectives, yet the core problems still persist in multimodal VAEs (Sutter et al., 2021, 2020). Due to these ongoing limitations, most multimodal studies have been constrained to low-dimensional datasets such as MNIST or PolyMNIST (LeCun and Cortes, 2010; Sutter et al., 2020). Additionally, Daunhawer et al. (2022) highlights the generative discrepancies in mixture-based models and raises concerns about the practicality of these models in real-world applications.
@@
On the other hand, diffusion models have emerged as a powerful class of generative models, gaining substantial attention in recent years for their ability to produce high-quality samples Ho et al. (2020); Dhariwal and Nichol (2021). These models operate by gradually transforming a data distribution into a noise distribution, with the goal of learning to reverse this process. The forward process is a Markovian sequence that incrementally adds noise to the data distribution over several timesteps, while the reverse process, typically modeled by a neural network, learns to remove the noise and reconstruct the data (Ho et al., 2020). A notable advantage of diffusion models is their capacity to generate exceptionally high-quality samples, achieving state-of-the-art performance in various generative tasks, including image synthesis and inpainting (Saharia et al., 2022; Lugmayr et al., 2022). Preechakul et al. (2022) introduced diffusion autoencoder, where the encoder resembles that of a VAE, but the decoder is modeled using a diffusion process. The key difference between a diffusion decoder and a traditional feed-forward decoder used in VAEs is that the diffusion decoder conditions the generation process on the latent representation rather than directly feeding it as input. This approach enables the diffusion autoencoder to learn meaningful representations as well as generate high-quality samples.
In this study, we introduce a novel multimodal generative model that combines the strengths of diffusion autoencoders and multimodal VAEs to learn a unified joint representation across multiple modalities. Traditional multimodal VAEs often face limitations in generating high-quality outputs from their joint representations, which may not capture all the details of all modalities  Palumbo et al. (2023); Hwang et al. (2021). These challenges are further exacerbated by the low-performance decoders typically used in standard VAEs, resulting in suboptimal model performance. To address these issues, we propose the use of a flexible diffusion decoder, which allows the compressed joint representation to effectively generate high-quality outputs for complex modalities while leveraging feedforward decoders for simpler modalities.
@@
To achieve this, we propose a multimodal ELBO that is integrated with diffusion decoders for modalities where feed-forward decoders have shown limitations, such as images, ensuring high-quality sample generation and improved representation capacity. Conversely, for modalities where standard decoders have proven effective, such as text or sparse outputs, we retain standard VAE decoders. This hybrid approach allows our model to harness the strengths of each generative model class while effectively addressing their respective limitations.
@@
Our model achieves better coherence in conditional generation among the generated modalities while generating high-quality samples. We use a mixture of product of experts (MoPoE) to fuse the representations from different modalities (Sutter et al., 2021). This approach enables our model to capture the correlations and dependencies between modalities. For unconditional generation tasks, where the goal is to generate samples from all modalities simultaneously without conditioning on any specific modality, we propose the use of an auxiliary model. This auxiliary model transforms a Gaussian noise distribution into the approximate posterior product of experts distribution. By leveraging this technique, our model gains increased flexibility and can perform well in both conditional and unconditional generation settings.
@@
Figure 2 illustrates the general design and architecture of our proposed model. The modality-specific encoders map the input data from different modalities into latent representations, which are then fused using the MoPoE mechanism. The fused representation is then passed to the respective decoders, which can be either diffusion decoder or a standard one, depending on the modality. Figure 2 shows how to perform unconditional sampling where the score-based model is employed to provide the necessary input distributions to the decoders.
@@
  • We introduce a novel generative multimodal autoencoder architecture that seamlessly integrates diffusion and standard decoders. This hybrid design allows our model to generate high-quality samples across modalities without compromising the coherence and consistency among the generated modalities. In doing so, our architecture strikes an optimal balance between sample quality and cross-modal coherence.
  • To enhance the model’s performance in unconditional generation settings, we propose to use a sampling technique that employs an auxiliary score-based model that transforms a Gaussian noise distribution into the approximate distribution of the PoE posterior to improve the model’s ability to generate coherent and diverse samples across all modalities in an unconditional manner for multimodal VAEs.
  • Our approach can be scaled to high-dimensional data modalities with good generative quality compared to previous approaches, which opens a new research direction toward multimodal VAEs.
Through these contributions, our model addresses the critical limitations of previous multimodal VAE approaches while retaining their desirable properties and extending their capabilities to handle high-dimensional data modalities effectively. Our approach paves the way for more powerful and flexible multimodal generative models capable of generating high-quality samples with enhanced coherence.
@@ 2 Related Works
Multimodal learning has emerged as a promising research direction in the deep learning field that has started from early works such as Ngiam et al. (2011); Tsai et al. (2019). Ngiam et al. (2011) demonstrated the potential of multimodal deep learning models in leveraging information from multiple modalities to improve performance on multiple tasks by learning combined features. As variational autoencoders (VAEs) gained traction for simultaneously learning representations and generation models within individual modalities, researchers soon explored extending the VAE framework to handle multiple modalities simultaneously, giving rise to multiple variants of multimodal VAEs.
@@
Pandey et al. (2022) proposed a two-stage training in which a unimodal VAE is trained in the first stage, and a diffusion model conditioned on the output of the VAE is trained in the second stage to refine the lower-quality output of a VAE. Similar approaches have also been proposed in multimodal works to refine the outputs of the image modalities of multimodal VAEs by  Palumbo et al. (2024); Wesego and Rooshenas (2024), but the diffusion model is not directly integrated to the ELBO like our proposed work.
Alternatively, Wesego and Rooshenas (2024); Bounoua et al. (2024) proposed to use unimodal VAEs to represent each modality and use an additional neural network to fuse the modalities instead of a product/mixture of experts using a two-step training process. They both select a score-based network to join the modalities in the latent space and sample from the joint distribution. Rombach et al. (2021) proposed text-to-image generation that employs a high-quality autoencoder to compress images into latent space, allowing for diffusion processes to occur in this reduced dimensionality. They integrated CLIP models using cross-attention to effectively represent and condition text inputs. Building upon this,  Xu et al. (2023); Tang et al. (2023) proposed multi-modal latent diffusion models that can generate any modality from any given condition. They expand Rombach et al. (2021) work to use multiple diffusion flows to achieve this. It’s worth noting that these advanced models require training on extremely large-scale datasets, which requires extensive computational resources.
@@ 3.1 Diffusion Autoencoder
A diffusion model can be formulated as a latent variable model where the marginal distribution is expressed as $p_{\theta}(\mathbf{x}_{0})=\int p_{\theta}(\mathbf{x}_{0:T})\mathrm{d}\mathbf{x}_{1:T}$ (Ho et al., 2020). The combined probability distribution $p_{\theta}(\mathbf{x}_{0:T})=p(\mathbf{x}_{T})\prod_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ is the reverse joint process where $p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{0},\mathbf{I})$. One can sample any intermediate time step $t$ in the forward process starting from the data point $\mathbf{x}_{0}$ using the Gaussian distribution $q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$ to add noise to the data. A neural network is trained to reverse this process where $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t))$. Ho et al. (2020) observed that using a simplified objective and predicting the noise is better and easier to implement. The objective is shown in eq. 1.
@@
A diffusion autoencoder combines the architectural setup of an autoencoder with a diffusion decoder (Preechakul et al., 2022). An encoder $q(\mathbf{z|x})$ projects the data into a latent representation $\mathbf{z}$. This latent representation is fed to the diffusion decoder which generates samples conditioned on $\mathbf{z}$. By training the model this way, the diffusion model can be used as a powerful representation where the encoder learns a meaningful latent representation of the data similar to an autoencoder (Hudson et al., 2023). The objective for a diffusion autoencoder is very similar to the diffusion objective, except we have additional conditioning from the encoder. It is trained by minimizing $\mathbb{E}_{\mathbf{x_{0},\epsilon_{t}}}[\left\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},\mathbf{z},t)\right\|^{2}]$ where $\mathbf{z}$ comes from the encoder network $q(\mathbf{z|x})$.
In addition, diffusion models can be trained to maximize the log-likelihood of training data by using a specific weight on the score-matching loss term (Song et al., 2021b). By using this objective, which is upper bounded by the log-likelihood, diffusion models can achieve log-likelihoods that are comparable to the best models. Specifically, with a stochastic differential equation (SDE) of the form $d\mathbf{x}=f(\mathbf{x},t)dt+g(t)d\mathbf{w}$ that diffuses data to noise with a drift coefficient $f$ and a diffusion coefficient $g$ where $\mathbf{w}$ is a standard Wiener process, we can use a weighting $\lambda(t)=g(t)^{2}$ for likelihood weighting. The denoising score-matching loss to train diffusion models for likelihood training is shown in eq.  2, where additional conditioning term $\mathbf{z}$ is added from the encoder to form the autoencoder setup and a Monte Carlo estimate is used for the expectation with $t$ sampled from uniform distribution, $\mathbf{x}$ from the training data, and $s_{\theta}()$ denotes the score function (Song et al., 2021b).
$L(\theta)=\,\mathbb{E}_{p(\mathbf{x})p(t)p(\mathbf{z|x})p_{t}(\mathbf{x}_{t}\mid\mathbf{x},\mathbf{z})}\Bigg{[}\frac{1}{2}\lambda(t)\Big{\|}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}\mid\mathbf{x},\mathbf{z})-s_{\theta}(\mathbf{x}_{t},\mathbf{z},t)\Big{\|}_{2}^{2}\Bigg{]}$ (2)
@@ 3.2 Multimodal VAE with Diffusion Decoder
Assuming that we have M modalities, the observed data is represented as $\mathbf{x}_{1:M}={\mathbf{x}^{(1)},\mathbf{x}^{(2)},\ldots,\mathbf{x}^{(M)}}$. Mixture-based multimodal models encode the posterior distribution using a mixture of experts of Gaussian distributions $q_{\phi}^{S}(\mathbf{z}|\mathbf{x})=\sum_{A\in\mathcal{S}}\omega_{A}q_{\phi}(\mathbf{z}|\mathbf{x}^{A})$ where $0\leq\omega_{A}\leq 1$ and $\mathcal{S}$ is a selected set of modalities formed from the powerset of the combination of modalities. Using a mixture of experts allows easy inference when some of the modalities are not observed, and training can be performed by sampling one component from the mixture at a time.  Daunhawer et al. (2022) highlighted that mixture-based multimodal models face an inherent limitation compared to unimodal VAEs due to the sub-sampling of modalities in the encoder. In these models, a restricted subset of modalities is used to encode information for all modalities, requiring the decoders to reconstruct all modalities from this limited input. This challenge is particularly evident in multimodal VAEs, where the decoder often struggles with the insufficient information provided by the encoder. However, incorporating a more powerful decoder can help address this issue effectively. For any subset $\mathcal{S}$ of modalities, a multimodal evidence lower bound (ELBO) can be derived for the log-likelihood of the multimodal data  Daunhawer et al. (2022):
@@
$\log p_{\theta}(\mathbf{x}_{1:M})\geq\sum_{A\in\mathcal{S}}\omega_{A}\Bigg{\{}\mathbb{E}_{q_{\phi}(\mathbf{z}\mid\mathbf{x}^{A})}\left[\log p_{\theta}(\mathbf{x}_{1:M}\mid\mathbf{z})\right]-D_{\text{KL}}\big{(}q_{\phi}(\mathbf{z}\mid\mathbf{x}^{A})\,\|\,p(\mathbf{z})\big{)}\Bigg{\}}.$ (3)
The full derivation is given in the Appendix  6.1 for reference. Daunhawer et al. (2022) demonstrates that this bound generalizes various multimodal VAE objectives, such as MoPoE Sutter et al. (2021), MMVAE Shi et al. (2019), and MVAE Wu and Goodman (2018).
Assuming that the modalities are conditionally independent given the latent representation $\mathbf{z}$ (i.e. $\mathbf{x}^{i}\perp\mathbf{x}^{\setminus i}\mid\mathbf{z}$), the ELBO can be written the following:
@@
$\log p(\mathbf{x}_{1:M})\geq\sum_{A\in\mathcal{S}}\omega_{A}\Bigg{\{}\mathbb{E}_{q_{\phi}(\mathbf{z}\mid\mathbf{x}_{A})}\left[\sum_{i=1}^{M}\log p_{\theta}(\mathbf{x}^{i}\mid\mathbf{z})\right]-D_{\text{KL}}\big{(}q_{\phi}(\mathbf{z}\mid\mathbf{x}^{A})\,\|\,p(\mathbf{z})\big{)}\Bigg{\}}.$ (4)
These conditional independencies enable us to support modality-specific decoders – specifically, feed-forward and diffusion-based decoders. To distinguish between them, we denote the number of modalities with feed-forward decoders as $M_{F}$ and those with diffusion-based decoders as $M_{D}$.
Proposition 1: The objective defined in eq.  3.2 is a valid lower bound on the marginal likelihood of the data under the proposed model.
@@
$\log$ $p(\mathbf{x}_{1:M})\geq\sum_{A\in\mathcal{S}}\omega_{A}\Bigg{\{}\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x}_{A})}\Bigg{[}\sum_{i=1}^{M_{F}}\log p_{\theta}(\mathbf{x}_{i}|\mathbf{z})+\sum_{j=1}^{M_{D}}\mathbb{E}_{\mathbf{x}_{jt}}\frac{1}{2}\lambda(t)\bigg{\|}\nabla_{\mathbf{x}_{jt}}\log p_{t}(\mathbf{x}_{jt}\mid\mathbf{x}_{j},\mathbf{z})$
$-s_{\theta}(\mathbf{x}_{jt},\mathbf{z},t)\bigg{\|}_{2}^{2}\Bigg{]}-D_{\text{KL}}\big{(}q_{\phi}(\mathbf{z}|\mathbf{x}_{A})\big{\|}p(\mathbf{z})\big{)}\Bigg{\}}.$ (5)
See Appendix  6.1 for the proof.